OpenStudy (anonymous):

The formula J=2.5sqrt{x}+27 models the percentage of full-time with jobs x years after 1980. Use the formula to project when 54.5% of full-time college students will have jobs.

OpenStudy (anonymous):

\[54.5=2.5\sqrt{x}+27\]

OpenStudy (anonymous):

\[81.5=2.5\sqrt{x}\]

OpenStudy (anonymous):

\[32.6=\sqrt{x}\]

OpenStudy (anonymous):

x=1063

OpenStudy (anonymous):

You wrote \[81.5=2.5\sqrt{x}\] If you check it again you'll notice you've added 27 instead of subtracting it.

OpenStudy (anonymous):

If you fix that the result will be x=121

OpenStudy (anonymous):

\[27.5=2.5\sqrt{x}\]

OpenStudy (anonymous):

\[11=\sqrt{x}\]

OpenStudy (anonymous):

x=121

OpenStudy (anonymous):

(y)

OpenStudy (anonymous):

in 2101, 54.5% students will have jobs?